### v2

# QKFormer: Hierarchical Spiking Transformer using Q-K Attention

Abstract
Spiking Transformers, which integrate Spiking Neural Networks (SNNs) with Transformer architectures, have attracted significant attention due to their potential for low energy consumption and high performance. However, there remains a substantial gap in performance between SNNs and Artificial Neural Networks (ANNs). To narrow this gap, we have developed QKFormer, a direct training spiking transformer with the following features: i) Linear complexity and high energy efficiency, the novel spike-form Q-K attention module efficiently models the token or channel attention through binary vectors and enables the construction of larger models. ii) Multi-scale spiking representation, achieved by a hierarchical structure with the different number of tokens across blocks. iii) Spiking Patch Embedding with Deformed Shortcut (SPEDS), enhances spiking information transmission and integration, thus improving overall performance. It is shown that QKFormer achieves significantly superior performance over existing state-of-the-art SNN models on various mainstream datasets. Notably, with comparable size to Spikformer (66.34 M, 74.81%), QKFormer (64.96 M) achieves a groundbreaking top-1 accuracy of 85.65% on ImageNet-1k, substantially outperforming Spikformer by 10.84%. To our best knowledge, this is the first time that directly training SNNs have exceeded 85% accuracy on ImageNet-1K. The code and models are available at https://github.com/zhouchenlin2096/QKFormer.
1 Introduction
Regarded as the third generation of neural networks [1], the brain-inspired Spiking Neural Networks (SNNs) are potential competitors to Artificial Neural Networks (ANNs) due to their high biological plausibility and high energy efficiency attributed to their event-driven properties [2]. Transformer, originally designed for natural language processing [3], has flourished in various computer vision tasks, including image classification [4, 5], object detection [6, 7, 8] and semantic segmentation [9, 10]. Spiking Transformers (Transformer-based SNNs) [11, 12, 13, 14, 15], which integrate spiking neural networks with transformer architecture, have attracted significant attention. This innovative combination provides great potential to develop advanced AI algorithms with high performance and low energy consumption.
As the architecture of the transformers is essential to the model’s performance [4, 5, 16, 8, 10], designing new architectures for transformer-based SNNs is quite challenging in terms of space requirements for the following reasons [11, 13, 15]. i). Spiking Self Attention (SSA) [11], the core module of spiking transformers, encodes Query, Key, and Value with sparse spikes. However, the computational complexity (especially space complexity) of SSA scales quadratically to the number of tokens (#tokens), and is the main obstacle to explore architecture that incorporate multi-level features. ii). SNNs process data across the time domain, necessitating a high level of computational and memory resources. This combination leads to considerable consumption of computational resources, making the training process highly demanding in terms of both memory and processing power.
To address these issues, we propose QKFormer with three innovations. i) Q-K attention with linear complexity and high energy efficiency. ii) A hierarchical architecture with decreasing number of tokens across blocks. iii) A novel patch embedding with deformed shortcut module. The linear complexity of Q-K attention is originated from the binary spike-form vector attention. This design lower the energy consumption and the space requirement. The hierarchical architecture starts from small patches and gradually merges neighboring patches in deeper spiking transformer layers with gradually decreasing #tokens, which enables multi-level spiking feature representation and benefits the model performance. The patch embedding with deformed shortcut facilitates spiking information transmission and integration. These merits make QKFormer achieve state-of-the-art performance in the SNN domain, in contrast to the previous transformer-based SNNs with spiking feature maps of a single resolution. Our main contributions are as follows:
1) We develop a novel spike-form Q-K attention mechanism, tailor-made for the spatio-temporal spiking patterns of SNNs, which can easily model the importance of token or channel dimensions with binary values. The Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($\mathbf{Q}$) and Key ($\mathbf{K}$).
2) We design a versatile and powerful Spiking Patch Embedding with Deformed Shortcut (SPEDS) module, which enhances spiking information transmission and integration thus improving the performance of spiking transformers significantly.
3) We build a direct-training hierarchical spiking transformer with different number of tokens across blocks, incorporating Q-K attention and SPEDS, named QKFormer. This marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs.
4) Extensive experiments show that the proposed model outperforms the state-of-the-art (SOTA) SNNs on various static and neuromorphic datasets. Notably, QKFormer has achieved a significant milestone, surpassing 85% top-1 accuracy on ImageNet with 4 time steps using the direct training approach for the first time.
Figure 1: Illustration of Q-K attention with the two versions of Q-K token attention (QKTA) and Q-K channel attention (QKCA). The inputs are binary spikes and there are only sparse additions and mask operations in Q-K attention. As a spike-driven module, Q-K attention efficiently models the token or channel attention through spike-form binary vectors, performing linear complexity to #tokens (or #channels) and high energy efficiency. Spiking Neuron (SN) in this work adopts the Leaky-Integrate-and-Fire (LIF) model, which is shown in Appendix. 6.1.
2 Related Work
Learning Methods of Spiking Neural Networks. At present, there are mainly two ways to obtain trained SNNs. One involves converting pre-trained ANNs to SNNs (ANN2SNN) [17, 18, 19, 20, 21, 22, 23], replacing the ReLU activation function in ANN with spiking neurons. However, This converted SNN suffers from long converting time steps and constraints on the original ANN design. Another method is to directly train SNNs[24], using surrogate gradient[25, 26, 27, 28] to address the non-differentiability of spike excitation function during backpropagation. The direct training method has received more attention due to its low latency and supporting flexible architectural exploration.
[…]
3 Method
3.1 Preliminary
Vanilla Self Attention. Vanilla self-attention (VSA) [3] in transformers has three floating-point key components: query ($\mathbf{Q}_{\mathcal{F}}$), key ($\mathbf{K}_{\mathcal{F}}$), value ($\mathbf{V}_{\mathcal{F}}$) which are calculated by learnable linear matrics and input $\mathbf{X}$. The calculation of VSA can be formulated as follows:
$\mathbf{Q}_{\mathcal{F}},\mathbf{K}_{\mathcal{F}},\mathbf{V}_{\mathcal{F}}=\mathbf{X}(\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}),$ (1)
$\operatorname{VSA}\left(\mathbf{Q}_{\mathcal{F}},\mathbf{K}_{\mathcal{F}},\mathbf{V}_{\mathcal{F}}\right)=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{\mathcal{F}}\mathbf{K}_{\mathcal{F}}^{\mathrm{T}}}{\sqrt{d}}\right)\mathbf{V}_{\mathcal{F}},$ (2)
where $\mathcal{F}$ denotes the floating-point form. Both floating-point matrix multiplication and softmax operation which contains exponent calculation and division, do not align with the properties of SNNs.
Spiking Self Attention. Spikformer [11] demonstrated a novel spike-form self-attention named Spiking Self Attention (SSA), using sparse spike-form $\mathbf{Q,K,V}$ without softmax operation and floating-point matrix multiplication. The calculation process of SSA is formulated as follows:
$\mathbf{I}=\mathrm{SN}_{I}\left(\mathrm{BN}_{I}\left(\mathbf{X}(\mathbf{W}_{I})\right)\right),\mathbf{I}\in(\mathbf{Q},\mathbf{K},\mathbf{V}),$ (3)
$\operatorname{SSA}^{\prime}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{SN}\left(\mathbf{Q}\mathbf{K}^{\mathrm{T}}\mathbf{V}*s\right),$ (4)
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathcal{R}^{T\times N\times D}$, the spike-form $\mathbf{Q},\mathbf{K},\mathbf{V}$ are computed by learnable linear layers. $s$ is a scaling factor. $\mathrm{SN}$ means spiking neuron layer. The calculation of SSA avoids floating-point multiplication, meeting the property of SNNs.
3.2 Q-K Attention
An overview of Q-K attention is shown in Figure 1. Both VSA and SSA use three key components and have $O(N^{2}d)$ or $O(Nd^{2})$ computational complexity, while our proposed Q-K Attention which has linear complexity and only uses two spike-form components: $\mathbf{Q}$ and $\mathbf{K}$, which are produced through learnable linear matrics.
$\mathbf{Q}=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(\mathbf{X}\mathbf{W}_{Q}\right)\right),\mathbf{K}=\mathrm{SN}_{K}\left(\mathrm{BN}\left(\mathbf{X}\mathbf{W}_{K}\right)\right),$ (5)
where $\mathbf{X}$ is the input spiking map. According to the detailed calculation mechanism of $\mathbf{Q},\mathbf{K}$, Q-K Attention can be divided into Q-K Token Attention (QKTA) and Q-K Channel Attention (QKCA).
Q-K Token Attention. We here assume $T=1$ and single head attention for mathematical description. After obtaining spike-form $\mathbf{Q},\mathbf{K}\in\mathcal{R}^{T\times N\times D}$, both $\mathbf{Q}$ and $\mathbf{K}$ can be formed as a spike matrix $N\times D$ ($N$ is the token number, $D$ is the channel number). QKTA process can be formulated as follows:
${\mathbf{A}}_{t}=\mathrm{SN}(\sum_{i=0}^{D}\mathbf{Q}_{i,j}),\quad\mathbf{X}^{\prime}={\mathbf{A}}_{t}\otimes\mathbf{K},$ (6)
where ${\mathbf{A}}_{t}$ is the $N*1$ token attention vector, which models the binary importance of different tokens. ${\mathbf{A}}_{t}$ is a spike-form vector, which is obtained by addition operations (row summation) of $\mathbf{Q}$ spike matrix and a following spiking neuron. $\otimes$ is the Hadamard product between spike tensors, which is equivalent to the mask operation. We apply the spike-form token attention vector ${\mathbf{A}}_{t}$ to the $\mathbf{K}$ spike matrix through the column mask operation (token mask), to obtain the output $\mathbf{X}^{\prime}$ of QKTA.
Q-K Channel Attention. The calculation process of Q-K channel attention is similar to the previous Q-K token attention, and can be formulated as : $\mathbf{A}_{c}=\mathrm{SN}(\sum_{j=0}^{N}\mathbf{Q}_{i,j}),\quad\mathbf{X}^{\prime}=\mathbf{A}_{c}\otimes\mathbf{K},$ where $\mathbf{A}_{c}$ is the $1*D$ channel attention vector, which models the binary importance of different channels. ${\mathbf{A}}_{t}$ is a spike-form vector, which is obtained by addition operations (column summation) of Q spike matrix and a following spiking neuron. Then, the output $\mathbf{X}^{\prime}$ of Q-K Channel Attention is obtained by the row mask operation (channel mask) between ${\mathbf{A}}_{t}$ and $\mathbf{K}$.
$\mathbf{X}^{\prime\prime}=\mathrm{SN}\left(\mathrm{BN}\left({\mathrm{Linear}}\left(\mathbf{X}^{\prime}\right)\right)\right).$ (7)
As shown in Formula.7, a post-linear layer is also required after obtaining $\mathbf{X}^{\prime}$ of Q-K Token or Channel Attention. In addition, the channel dimension is $D/h$ in the multi-head Q-K attention, where $h$ is the head number. In this work, the spiking neuron uses the LIF model [28]. Same with [11], time step $T$ is an independent dimension for the spiking neuron layer. In other layers, it is merged with the batch size. We exploit QKTA in our experiments by default.
Linear Computational Complexity of Q-K Attention. As shown in Table 1, the time complexity of Q-K attention varies based on the implementation approach. Specifically, when utilizing spike-form broadcasted element-wise multiplication, $\otimes$, the time complexity can reach up to $O(N*D)$. When applying mask operation, the time complexity of Q-K attention is only $O(N)$ or $O(D)$. The space complexity of Q-K attention with the whole process is $O(N*D)$ at most, which is caused by the self-storage consumption Q and K matrix. In terms of the space complexity of attention operation, Q-K attention only requires an extra $1*D$ or $N*1$ space to store the attention vector with the space complexity of $O(N)$ or $O(D)$. The linear complexity of Q-K attention makes it possible to successfully explore the large-scale hierarchical architecture SNN model.
High Energy Efficiency of Q-K Attention. As a spike-driven attention module, the linear multiplication is transformed into sparse addition. Mask operation can be implemented on neuromorphic chips through addressing algorithms [29] or AND logic operations[30] with negligible power consumption. Compared with SSA, Q-K attention is much more energy-efficient, which comes from the following reasons: i) Q-K attention only adopts two spike-form components for spike [0, 1] operation without the V input and thus has less synaptic computing. ii) Q-K attention has much fewer spiking matrix operations due to its linear complexity of $O(N)$ or $O(D)$. iii) Q-K attention discards the scale operation of SSA, which leads to reduced power consumption further.
Table 1: Computational complexity comparison. $N$ is the token number, $D$ is the channel number.
Methods VSA [3] SSA [11] SDSA [13] QKTA QKCA
Time complexity $O(N^{2}D)$ $O(N^{2}D)$ $O(ND)$ $O(D)$ $O(N)$
Space complexity $O(N^{2}+ND)$ $O(N^{2}+ND)$ $O(ND)$ $O(N)$ $O(D)$
3.3 No Scaling Factors in Q-K Attention
In VSA [3], assume that $\mathbf{q}_{i}\left(\mathbf{q}_{i}\in R^{1\times d},\mathbf{Q}\in R^{m\times d}\right)$ and $\mathbf{k}_{i}\left(\mathbf{k}_{i}\in R^{1\times d},\mathbf{K}\in R^{m\times d}\right)$ are independent random variables with a mean of 0 and a variance of 1, then each element in the product of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}$ has mean 0 and variance $d$. The variance magnitude of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}$ grows with the embedding dimension $d$, which can result in gradient vanishing issues after softmax operation. Therefore, The product of matrices $\mathbf{Q}$ and $\mathbf{K}$ in VSA [3] is scaled by a factor $\frac{1}{\sqrt{d}}$ in Eq. 2 to normalize the product to variance 1. Though the softmax function is not adopted due to its non-spike operations (division, exponential operation) in SNNs, SSA-based [11] SNNs will suffer obvious performance degradation even cannot converge without scaling because the variance of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}\mathbf{V}$ output is too large (Assuming that all the spiking elements are independent random variables and subject to Bernoulli Distribution). However, Q-K attention can discard scaling operations thus reducing power consumption because the variance of Q-K attention is much smaller than SSA (e.g. the max theoretical variance of Q-K token attention is only about 1 / 200 of SSA). The detailed analysis can be found in the Appendix.6.2 and Section.4.3.
3.4 QKFormer
As the computational complexity (especially space complexity) of SSA is quadratic to #tokens, previous direct training spiking transformers are all limited to straight-through structures. Combining SSA with hierarchical architecture directly will lead to memory explosion easily when training spiking transformers. To overcome these issues, we proposed a hierarchical spiking transformer based on Q-K attention, named QKFormer, which constructs hierarchical spiking feature maps with linear computational complexity to #tokens or #channels.
Figure 2: The overview of QKFormer, a hierarchical spiking transformer with Q-K attention.
Overall Hierarchical Architecture. The overview of QKFormer is presented in Figure 2. The input form can be formulated as $(T_{0}\times H\times W\times n)$. In static RGB image datasets, $T_{0}=1$ and $n=3$. In temporal neuromorphic datasets, the input $T_{0}=T$, while $n=2$. In our implementation, we use a patch size of $4\times 4$ and thus the input feature dimension ($4\times 4\times n$) of each patch is projected into a spike-form arbitrary dimension (denoted as $C$) in Spiking Patch Embedding with Deformed Shortcut 1 (SPEDS-1), which together with the following QKFormer blocks are referred to as "Stage 1". The number of tokens in Satge 1 is $(\frac{H}{4}\times\frac{W}{4})$. To produce a hierarchical spiking representation, the number of tokens is reduced in SPEDS-2 and SPEDS-3 as the network goes deeper. Both SPEDS-2 and SPEDS-3 reduce the number of tokens by a patch size of $2\times 2$ (2×2$\times$2 × downsampling of resolution), and the number of channels is transformed into $2C$ and $4C$, respectively. We denote the SPEDS-2 and the following QKFormer blocks as "Stage 2", which reduces the number of tokens $(\frac{H}{8}\times\frac{W}{8})$. While SPEDS-3 and the following Spikformer or QKormer blocks are referred to as "Stage 3" with $(\frac{H}{16}\times\frac{W}{16})$ tokens. The number of spiking transformer blocks (QKFormer or Spikformer) in each stage are $N_{1}$, $N_{2}$, and $N_{3}$, respectively. These stages jointly produce a hierarchical spike-form representation.
Mixed Spiking Attention Integration. In the former stage of a hierarchical architecture model, the number of channels is small while the number of tokens is large. In the latter stage, the channel number is large while the token number is small. Thus it leads to suboptimal performance when we only use a single type of Q-K attention in a hierarchical architecture model. Therefore, we use mixed spiking attention integration in QKFormer. QKTA is conducted in the former stage in hierarchical architecture, and we could choose QKCA or SSA in the latter stage. In the subsequent experiments, we use SSA in the last stage of QKFormer and QKTA in the former stages by default.
QKFormer Blocks. Similar to the standard transformer encoder block, a QKFormer block contains a Q-K Attention module (QKTA or QKCA) and a Spiking MLP (SMLP) block, which can be formulated as follows:
$\mathbf{X}_{l}^{\prime}=\operatorname{QKTA}\left(\mathbf{X}_{l-1}\right)+\mathbf{X}_{l-1},\mathbf{X}_{l}^{\prime}\in R^{T\times N\times D},$ (8)
$\mathbf{X}_{l}=\operatorname{SMLP}\left(\mathbf{X}_{l}^{\prime}\right)+\mathbf{X}_{l}^{\prime},\mathbf{X}_{l}\in R^{T\times N\times D}.$ (9)
At last, a fully connected layer is used as the classifier behind the last block.
3.5 Spiking Patch Embedding with Deformed Shortcut.
Residual shortcuts in SNNs [28] can implement identity mapping, which reduces information loss (facilitates information transmission and integration) in spike communication, thus ensuring the network can be well-behaved in a depth-insensitive way. Previous spiking transformers [11, 12, 13] use the residual shortcuts to achieve identity mapping, mainly focusing on the spiking attention block and spiking MLP block, and lacking identity mapping in patch embedding across the downsampling block. The input and output of a spiking patch embedding block in QKFormer have different channel and token numbers. To realize residual learning in spiking patch embedding, we can perform a lightweight linear projection $\mathbf{W}_{d}$ in the shortcut connections to match the channel and token numbers, thus realizing the identity mapping cross downsampling blocks in spiking patch embedding. Given the input spiking map $\mathbf{X}$, the process of patch embedding can be formulated as follows:
$\mathbf{Y}=\mathcal{F}\left(\mathbf{X},\left\{\mathbf{W}_{i}\right\}\right)+\mathrm{SN}(\mathbf{W}_{d}\mathbf{X}).$ (10)
[…]
There are mainly two types of residual shortcuts in deep SNNs. Formula.10 shows the patch embedding in the way of activation-before-addition [28, 11]. The other way of the patch embedding with the pre-activation residual shortcut [31, 12, 13] can be formulated as follows:
$\mathbf{Y}=\mathrm{SN}(\mathcal{G}\left(\mathbf{X},\left\{\mathbf{W}_{j}\right\}\right)+\mathbf{W}_{d}\mathbf{X}),$ (11)
[…]
Table 2: Results on ImageNet-1K. Power is calculated as the average theoretical energy consumption when predicting an image from ImageNet test set. The power data for QKFormer and ANNs is evaluated according to Appendix.6.6, and the power data for other works were obtained from related papers. "A2S" denotes "ANN-to-SNN", "HST-$L$-$D$" denotes "Hierarchical Spiking Transformer" with $L$ encoder blocks and $D$ channels. HST-10-768 and HST-10-768∗∗ means HST-10-768 with 2882 and 3842 input size for inference. The top-5 accuracy of QKFormer (HST-10-768∗∗) is 97.74%.
[…]
RMP[21] A2S VGG-16 2242 39.90 - 2048 73.09
QCFS[22] A2S ResNet-18 2242 11.70 - 1024 74.32
MST[23] A2S Swin Transformer-T 2242 28.50 - 512 78.51
SEW ResNet[28] SNN SEW-ResNet-34 2242 21.79 4.89 4 67.04
SNN SEW-ResNet-101 2242 44.55 8.91 4 68.76
SNN SEW-ResNet-152 2242 60.19 12.89 4 69.26
Spikformer[11] SNN Spikformer-8-384 2242 16.81 7.73 4 70.24
SNN Spikformer-8-512 2242 29.68 11.58 4 73.38
SNN Spikformer-8-768 2242 66.34 21.48 4 74.81
Spikingformer[12] SNN Spikingformer-8-384 2242 16.81 4.69 4 72.45
SNN Spikingformer-8-512 2242 29.68 7.46 4 74.79
SNN Spikingformer-8-768 2242 66.34 13.68 4 75.85
S-Transformer[13] SNN S-Transformer-8-384 2242 16.81 3.90 4 72.28
SNN S-Transformer-8-512 2242 29.68 1.13 1 71.68
SNN S-Transformer-8-512 2242 29.68 4.50 4 74.57
SNN S-Transformer-8-768 2882 66.34 6.09 4 77.07
ViT[4] ANN ViT-B/16 3842 86.59 254.84 1 77.90
DeiT[32] ANN DeiT-B 2242 86.59 80.50 1 81.80
ANN DeiT-B 3842 86.59 254.84 1 83.10
Swin[8] ANN Swin Transformer-B 2242 87.77 70.84 1 83.50
ANN Swin Transformer-B 3842 87.77 216.20 1 84.50
QKFormer SNN HST-10-384 2242 16.47 15.13 4 78.80
SNN HST-10-512 2242 29.08 21.99 4 82.04
SNN HST-10-768 2242 64.96 8.52 1 81.69
SNN HST-10-768 2242 64.96 38.91 4 84.22
SNN HST-10-768 2882 64.96 64.27 4 85.25
SNN HST-10-768∗∗ 3842 64.96 113.64 4 85.65
4 Experiments
4.1 Results on ImageNet-1k Classification
Experimental Setup on ImageNet. In this experiment, we use AdamW as the optimizer, which is adopted with a base learning rate of $6\times 10^{-4}$. The actual learning rate was calculated as BatchSize/256 multiplied by the base learning rate. The batch size is set to $512$, which is realized by accumulated gradient iterations [33] and distributed across 8 Nvidia V100 GPUs. We trained QKFormer for $200$ epochs. In addition, following DeiT [32], data augmentation techniques including RandAugment [34], random erasing [35], and stochastic depth [36] are employed in this study. The number of blocks in the three stages is set as {1, 2, 7} respectively.
Main Results on ImageNet. The experimental results demonstrate the superior performance of our proposed QKFormer, surpassing previous works’ performance by a large margin (Table 2). QKFormer (64.96 M) achieves 85.65% top-1 accuracy and 97.74% top-5 accuracy on ImageNet. To begin with, we compare our model with the baseline spiking transformer (i.e., Spikformer [11]). Our QKFormer models have slightly fewer parameters but much higher performance. For example, our QKFormer (64.96 M, 85.65%) significantly outperforms Spikformer (66.34 M, 74.81%) by 10.84%. In addition, compared with SDSA, our Q-K attention has lower computational complexity (shown in Table 1) and our QKFormer has much higher performance than S-Transformer (built by SDSA) [13]. In detail, QKFormer outperforms S-Transformer by 7.55%, 7.47%, and 8.58% respectively on three models with comparable #parameters. Finally, Our QKFormer outperforms the SOTA ANN-to-SNN model MST [23] by 7.14% and has much fewer time steps meanwhile. To our best knowledge, this is the first time that a direct training SNN model has achieved an accuracy of over 85% on ImageNet-1k.
Comparing with ANN Models on ImageNet. Our QKFormer is an event-driven SNN model, whose output is in binary form (either 0 or 1), the multiplications of activations and weights can be transformed into sparse addition, thus enjoying high energy efficiency. It should be noted that hierarchical architecture will lead to the power increment of QKFormer. This is still very cost-effective compared with ANN models. For instance, QKFormer (64.96M, 85.65%, SNN, 113.64mJ) Vs. Swin Transformer (88M, 84.5%, ANN, 216.20mJ) [8] Vs. DeiT-B (86M, 83.1%, ANN, 254.84mJ) [32] Vs. ViT (85.59M, 77.9%, ANN, 254.84mJ). [4]. Under the same experiment conditions without pre-training or extra training data, our QKFormer has surpassed the most well-known Transformer-based ANNs in performance while maintaining high energy efficiency.
Table 3: Comparision on CIFAR10, CIFAR100, DVS128, CIFAR10-DVS. "Param" denotes "Parameter (M)", "Acc" denotes "Top-1 Accuracy (%)", "$T$" denotes "Time Step".
Method CIFAR10 CIFAR100 DVS128 CIFAR10-DVS
Param $T$ Acc Param $T$ Acc Param $T$ Acc Param $T$ Acc
Spikformer [11] 9.32 4 95.51 9.32 4 78.21 2.57 16 98.3 2.57 16 80.9
Spikingformer [12] 9.32 4 95.81 9.32 4 78.21 2.57 16 98.3 2.57 16 81.3
CML [14] 9.32 4 96.04 9.32 4 80.02 2.57 16 98.6 2.57 16 80.9
S-Transformer[13] 10.28 4 95.60 10.28 4 78.4 2.57 16 99.3 2.57 16 80.0
STSA[15] -- -- -- -- -- -- 1.99 16 98.7 1.99 16 79.93
ResNet-19 (ANN) 12.63 1 94.97 12.63 1 75.35 -- -- -- -- -- --
Trasnformer (ANN) 9.32 1 96.73 9.32 1 81.02 -- -- -- -- -- --
QKFormer 6.74 4 96.18 6.74 4 81.15 1.50 16 98.6 1.50 16 84.0
4.2 Results on CIFAR and Neuromorphic Datasets
CIFAR Classification. In this experiment, the QKFormer is trained for 400 epochs with a batch size of 64 following previous works: Spikformer [11], Spikingformer [12]. Following Spikformer, we use 4 blocks in QKFormer in total, which are distributed {1, 1, 2} in three stages. Due to the hierarchical architecture design, our QKFormer model has only 6.74 M parameters in this case. The results on CIFAR datasets are shown in Table 3. For CIFAR10, our model achieved 96.18% accuracy with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 0.67% and reduces 2.58 M parameters meanwhile. For CIFAR100, our model achieved 81.15% with 6.74 M parameters. Our proposed QKFormer outperforms Spikformer by 2.94% and reduces 2.58 M parameters meanwhile.
Neuromorphic Classification. We compare our method with SOTA methods on both CIFAR10-DVS and DVS-Gesture datasets. In this experiment, We utilize a mini QKFormer model with 1.50 M parameter, which has {0, 1, 1} blocks in three stages. The max patch embedding dimension is set to 256. The training process involves 200 epochs for DVS128 Gesture and 106 epochs for CIFAR10-DVS. The number of time steps of the spiking neuron is 10 or 16. The experimental results of temporal neuromorphic classification are presented in Table 3. For DVS128-Gesture dataset, our model with 1.50 M parameters achieves 98.6% accuracy using 16 time steps and 98.3% accuracy using 10 time steps. For CIFAR10-DVS dataset, our model achieves 84.0% accuracy with only 1.50 M parameters using 16 time steps. Our proposed QKFormer significantly outperforms Spikformer by 3.1% while reducing 1.07 M parameters. In addition, our model with 10 time steps achieves 83.8% accuracy, which outperforms Spikformer by 4.9% and outperforms the SOTA model (Spikingformer) by 3.9%.
Figure 3: The visualization and memory consumption of QKTA. 3 is the visualization of Q-K token attention. The white dot means value 1, while the black one means value 0. 3 shows the comparison of memory costs between QKTA and SSA under different token numbers. $N$ is the token number.
Figure 4: (a) shows the variance and expectation of SSA, (b) shows the variance and expectation of QKTA. Assume that all the spike elements (either 0 or 1) in SSA and QKTA are independent random variables and subject to Bernoulli distribution.
4.3 Analyses on Q-K Attention
Attention Visualization. In this part, we visualize the Q-K token attention (Stage 1 and Stage 2 of the QKFormer model) on ImageNet. As shown in Figure 3, ${\mathbf{A}}_{t}$ is the $N*1$ token attention vector, and $\mathbf{X}^{\prime}$ is the output of the attention process, which is obtained by the mask operation between matrix $\mathbf{K}$ and attention vector ${\mathbf{A}}_{t}$. Specifically, the longitudinal axis denotes the channel index of one head, while the horizontal axis denotes the token index. The #tokens in stage 1 and stage 2 are $56^{2}$ and $28^{2}$, respectively. To facilitate visualization, we choose a continuous segment with a length of 100 extracted from the whole token vector. The visualization shows Q-K attention can lead to high sparsity of spikes.
Memory Consumption. In this experiment, we compare the memory consumption between QKTA (Formula.6) and SSA (Formula.4) under different token numbers. We calculate the memory consumption of a QKTA and an SSA on a GPU by forwarding the input tensor $(T,B,C,N)$. To facilitate the statistics of the impact of #tokens $N$ on memory consumption, the #channels $C$ is set to 256, and the time step $T$ and batch size $B$ are set to 1. The experiment result is shown in Figure 3. With the increment of #Tokens, SSA consumes much more GPU memory than QKTA, of which the complexity is linear to #Tokens. For example, SSA consumes about 10×10$\times$10 × GPU memory than QKTA when $\sqrt{N}=50$.
Table 4: Spike firing rates in QKFormer blocks.
QKFormer Block Stage1 (fr) Stage2 (fr)
QKTA $\mathbf{Q}$ 0.0432 0.0231
$\mathbf{K}$ 0.1784 0.0847
${\mathbf{A}}_{t}$ 0.3477 0.2655
$\mathbf{X}^{\prime}$ 0.0832 0.0350
$\mathbf{X}^{\prime\prime}$ 0.1478 0.0577
SMLP Layer1 0.0518 0.0246
Layer2 0.2733 0.1869
Spike Firing Rates in QKFormer Blocks. In this experiment, we calculate the spike firing rates of QKFormer blocks of the trained QKFormer (64.9M) on the ImageNet-1K test set with the 224 $\times$ 224 input resolution. The average spike firing rates of the QKFormer blocks in Stage1 and Stage2 are shown in Table 4. Note that the spike-form $\mathbf{X}^{\prime}$ is obtained by column mask operation (token mask) between ${\mathbf{A}}_{t}$ and $\mathbf{K}$. In fact, the summation operation in the Q-K attention causes $\mathbf{Q}$ to become significantly sparser compared to $K$ when the network converges. Specifically, $\mathbf{Q}$ in stage 1 has a fire rate of 0.0432, while $\mathbf{K}$ has 0.1784. After the accumulation operation along $D/h$ of the multi-head QKTA version, the LIF neuron (${\mathbf{A}}_{t}$) exhibits a typical averaged fire rate of 0.3477.
The Variance and Expectation of QKTA. Figure.4 visualize the variance and expectation of QKTA (Formula.15 and 16 in Appendix.6.2) and SSA (Formula.19 and 20 in Appendix.6.2). $N$ is set as 196 and $d$ is set as 64, respectively. We can find SSA has a much larger variance and expectation than QKTA on the whole. For example, the maximum theoretical variance of QKTA is 16, but the maximum theoretical variance of SSA is over 3000. This is the main reason that Q-K attention can discard scaling operations thus reducing power consumption, but SSA can not.
4.4 Ablation Study
SPEDS Module. In this experiment, We replaced the Spiking Patch Splitting (SPS) module in Spikformer with Spiking Patch Embedding with Deformed Shortcut (SPEDS) module, while other conditions remain unchanged. The results (Table 5) show that the SPEDS module is essential to QKFormer on both static and neuromorphic datasets. In addition, the addition of SPEDS to Spikformer leads to great gains in CIFAR100 (+2.05%) and CIFAR10-DVS (+1.30%), which further verified the effectiveness of SPEDS.
Table 5: Ablation studies of SPEDS module.
Model CIFAR100 (Acc) CIFAR10-DVS (Acc)
QKFormer (QKTA + SSA, baseline) 81.15% 84.00%
QKFormer (QKTA + SSA, w/o SPEDS) 80.08% 83.40%
Spikformer (SSA, w/o scaling) 76.95% 79.30%
Spikformer (SSA) 78.21% 80.90%
Spikformer (SSA) + SPEDS 80.26% 82.20%
Table 6: Ablation studies of Q-K Attention.
Model CIFAR100 (Acc, Param) CIFAR10-DVS (Acc, Param)
QKFormer (QKTA + SSA, baseline) 81.15%, 6.74M 84.00%, 1.50M
QKFormer (QKCA + SSA) 81.07%, 6.74M 84.30%, 1.50M
QKFormer (QKTA + QKCA) 81.04%, 6.44M 83.10%, 1.44M
QKFormer (SSA) 81.23%, 6.79M 84.10%, 1.52M
QKFormer (QKCA) 81.00%, 6.44M 80.70%, 1.44M
QKFormer (QKTA) 79.09%, 6.44M 80.70%, 1.44M
Mixed Spiking Attention Integration with Q-K Attention. In this part, we show different integration strategies of QKCA, QKTA, and SSA. The baseline is our QKFormer (QKTA + SSA, 6.70M). The experimental results (Table 6) show that using a single type of Q-K attention (QKTA or QKCA only) in a hierarchical architecture model leads to suboptimal performance. In particular, the performance decline in QKTA is more obvious. While the mixed spiking attention solutions, such as QKFormer(QKTA + QKCA), QKFormer(QKTA + SSA), and QKFormer(QKCA + SSA) can achieve comparable performance to QKFormer(SSA) while requiring fewer parameters and much fewer memory resources (Figure 3). Consequently, the mixed spiking attention solutions are well-suited for larger architectures and more challenging scenarios when considering both computational efficiency and performance.
Table 7: Ablation studies of RC, SN, TS.
Model CIFAR100 (Acc)
QKFormer (baseline) 81.15%
QKFormer (ABA->PA) 81.18%
QKFormer (LIF->IF) 80.95%
QKFormer (LIF->PLIF) 81.12%
QKFormer (T=1) 78.51%
QKFormer (T=2) 80.08%
QKFormer (T=4) 81.15%
QKFormer (T=6) 81.30%
Residual Connection (RC) & Spiking Neuron (SN) & Time Step (TS). The experimental results are shown in Table 7. In this block, we replaced the Activation-Before-Addition (ABA) [28, 11] residual connection of QKFormer with the Pre-activation (PA) [31, 37] way, and the performance slightly improved. In addition, we replaced the LIF spiking neuron with Integrate-and-Fire (IF) and Parametric-Leaky-Integrate-and-Fire (PLIF) [38], which led to slight performance degradation. The accuracy regarding different simulation time steps of QKFormer is shown in the last column. When we increase the simulation time steps, the performance of QKFormer can be further improved. Specifically, QKFormer achieves 81.30 % accuracy on CIFAR 100 when T=6.
5 Conclusion
In this work, we design a novel spike-form Q-K attention considering the properties of SNNs, which can easily model the importance of token or channel dimensions through binary vectors. Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($\mathbf{Q}$) and Key ($\mathbf{K}$). We design a versatile and powerful Spiking Patch Embedding with Deformed Shortcut (SPEDS), which enhances spiking information transmission and integration, thus improving the performance of spiking transformers. In addition, we develop a hierarchical spiking transformer based on the proposed Q-K attention and SPEDS in a direct training way, named QKFormer, which marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs. Extensive experiments show that the proposed model achieves SOTA performance on both static and neuromorphic datasets. Note that QKFormer achieved top-1 accuracy beyond 85% on ImageNet-1k with 4 time steps using the direct training way for the first time. With its powerful performance, we aim for our investigations to instill optimism in the application of SNNs.
Limitation. Currently, our model is limited to image / DVS classification tasks. We will extend this work to more tasks, such as segmentation, detection, and language tasks, to test the generalizability in the further. In addition, we will explore efficient and high-performance network architectures with fewer time steps based on Q-K attention and other efficient modules, to further reduce the training consumption.